A 40-ft ladder is leaning against a wall. If the ladder reaches 20 ft up the wall,how far away from the base of the wall is the foot of the ladder?
Answer by rfer(16322)

a^2+20^2=40^2
a^2+400=1600
a^2=1600-400
a^2=1200
sqrt1200=34.64 ft